# Crowdsourcing Ubiquitous Indoor Localization with Non-Cooperative Wi-Fi Ranging

### v1

@@ 6.4. Comparison with Baselines
Core Location (iOS) 13.33 13.77 5.03
@@
Fused Location Provider (Android) + OpenLocate 7.71 6.74 3.98
@@
We show the output of PeepLoc s AP localization in Fig. 10 and PeepLoc s client localization in Fig. 11.
@@ 7. Discussion
8. Related Work
Fingerprinting. Fingerprinting remains one of the most widely adopted techniques for indoor localization. It leverages ambient signal measurements from various sensors (e.g. Wi-Fi, BLE, magnetometer) to create a map of the environment. The seminal RADAR system (bahl2000radar, ) demonstrated that Wi-Fi RSS fingerprints could localize users with median errors of 2–3 meters in an office environment without additional hardware infrastructure. Subsequent work has improved on this framework using better statistical models e.g. Horus (youssef2005horus, ), including more sensors (wang_no_2012, ), and incorporating deep learning (nowicki2017wifi, ). Fingerprinting has resulted in many commercial-grade crowdsourced systems (e.g. Apple’s indoor maps program (apple_indoor_maps, ) and Tencent’s indoor localization system (ni_experience_2022, )). Despite their popularity, fingerprinting systems require costly and time-consuming data collection, and their performance degrades in dynamic environments.
Two-Way Wi-Fi Ranging. To overcome the setup overhead of fingerprinting, two-way time-of-flight (ToF) based methods have been explored. The IEEE 802.11mc standard, also known as Fine Time Measurement (FTM), enables devices to estimate round-trip propagation delay between themselves and Wi-Fi APs, allowing direct ranging without prior training data. Ibrahim et al. evaluated the feasibility of Wi-Fi RTT on commercial smartphones and achieve sub-2 meter accuracy in line-of-sight indoor settings (ibrahim2018rtt, ). Although two-way Wi-Fi ranging has high accuracy, it is supported only by a few devices and AP manufacturers on the market (e.g. Aruba Open Locate (troymart_its_2023, )), limiting its widespread deployment.
One-Way Wi-Fi Ranging With Channel. Unlike two-way ranging methods, one-way ranging does not require cooperation from the Wi-Fi APs. One class of localization systems relying on one-way ranging methods infer location from observing the wireless channel as measured at the device. These include SpotFi (kotaru2015spotfi, ), which uses AoA and ToF extracted from commodity Wi-Fi chipsets using CSI and MUSIC algorithms, and Chronos (vasisht2016chronos, ) which measures packet arrival times across multiple frequency bands to estimate distance between a device and AP pair. These methods require access to low-level PHY data and customized firmware, limiting their practical deployment.
Non-Cooperative Wi-Fi Ranging. Another class of systems uses one-way ranging measurements without relying on the channel information. These include our system, Wi-Peep (abedi_non-cooperative_2022, ), and Horn’s system (horn_indoor_2022, ). Wi-Peep tackles the problem of localizing devices inside an indoor environment using a GPS-equipped drone flying around the building. Since the drone functions as the anchor point for localization in this case, it is not suitable as a permanently deployed localization system. The system proposed by Horn (horn_indoor_2022, ) uses non-cooperative Wi-Fi ToF as an indoor localization system, but assumes the locations of the APs within the environment are known beforehand. By contrast, our system is capable of automatically locating where the APs are inside an environment through crowdsourcing.
@@ 9. Conclusion
We present PeepLoc, a Wi-Fi ranging-based indoor positioning system that can be ubiquitously deployed. PeepLoc requires no additional hardware beyond that present in commodity smartphones, and no infrastructure requirements beyond pre-existing Wi-Fi APs. Unlike existing indoor localization systems, PeepLoc requires no cooperation from the infrastructure, and can bootstrap itself from pedestrian data with no dedicated human effort. In this paper, we show how PeepLoc can be deployed in various campus buildings, showing improved performance over both fingerprinting-based approaches as well as state-of-the-art Wi-Fi ranging-based approaches (e.g. FTM). Future work can extend PeepLoc to work on buildings with multiple floors, or handle changes in the arrangement of Wi-Fi APs.

### v2

@@ 6.3. Comparison with State-of-the-Art
Core Location (iOS) (apple:corelocation, ) 13.33 13.77 5.03
@@
Fused Location Provider (Android) (google:fusedlocationprovider, ) + OpenLocate (troymart_its_2023, ) 7.71 6.74 3.98
@@
6.4. Ranging Model Ablation Study
We conduct an ablation study on the per-AP ranging model by comparing the end-to-end client localization accuracy with and without the per-AP ranging model enabled (in which case we use a fixed slope of $\frac{c}{2}$). We summarize the results in Fig. 9(c). As the graph shows, the per-AP ranging model is crucial for accurate location estimation.
6.5. Visualization
For completeness, we show the output of PeepLoc’s AP localization algorithm in Fig. 10 and PeepLoc’s client localization algorithm in Fig. 11 respectively. For ease of reference, the results are superimposed onto the building floorplans.
@@ 8. Conclusion
We present PeepLoc, a Wi-Fi ranging-based indoor positioning system that can be ubiquitously deployed. PeepLoc requires no additional hardware beyond that present in commodity smartphones, and no infrastructure requirements beyond pre-existing Wi-Fi APs. Unlike existing indoor localization systems, PeepLoc requires no cooperation from the infrastructure, and can bootstrap itself from pedestrian data with no dedicated human effort. In this paper, we show how PeepLoc can be deployed in various campus buildings, showing improved performance over both fingerprinting-based approaches as well as state-of-the-art Wi-Fi ranging-based approaches (e.g. FTM). Future work can extend PeepLoc to work on buildings with multiple floors, or handle changes in the arrangement of Wi-Fi APs over time.